# Emergent Cooperation for Energy-efficient Connectivity via Wireless Power Transfer

### v1

@@ Abstract
This paper addresses the challenge of incentivizing energy-constrained, non-cooperative user equipment (UE) to serve as cooperative relays. We consider a source UE with a non-line-of-sight channel to an access point (AP), where direct communication may be infeasible or may necessitate a substantial transmit power. Other UEs in the vicinity are viewed as relay candidates, and our aim is to enable energy-efficient connectivity for the source, while accounting for the self-interested behavior and private channel state information of these candidates, by allowing the source to “pay” the candidates via wireless power transfer (WPT). We propose a cooperation-inducing protocol, inspired by Myerson auction theory, which ensures that candidates truthfully report power requirements while minimizing the expected power used by the source. Through rigorous analysis, we establish the regularity of valuations for lognormal fading channels, which allows for the efficient determination of the optimal source transmit power. Extensive simulation experiments, employing real-world communication and WPT parameters, validate our theoretical framework. Our results demonstrate a 91% reduction in outage probability with as few as 4 relay candidates, compared to the non-cooperative scenario, and as much as 48% source power savings compared to a baseline approach, highlighting the efficacy of our proposed methodology.
@@ I Introduction
The forthcoming 6G wireless standard promises to revolutionize wireless communication by delivering significant improvements in reliability, data rates, and energy efficiency [1]. Achieving the desired technical performance depends on a number of emerging technologies and paradigms, which in turn bring their own challenges. For instance, high frequency communication (e.g., mmWave, THz) enables vastly increased capacity, but suffers from acute shadowing and penetration losses [2]. At the same time, given the ever-increasing number of connected devices, there is a growing need for scalable system design, with reduced reliance on centralized coordination mechanisms, to produce the desired performance.
This paper proposes the use of wireless power transfer (WPT) [3] to induce cooperative behavior among non-cooperative communication devices, with the end goal of enabling connectivity and/or reducing energy consumption. More specifically, we consider a setting in which a large scale blockage (e.g., a wall) obstructs the line-of-sight (LoS) link between a user equipment (UE) and an access point (AP), so that direct communication is challenging. To reduce the power required to connect to the AP, the blocked user may offer to send a payment of power (via WPT) to a more advantageously placed UE, in exchange for relay services. The core problem we address in this paper is the identification of the optimal relay candidate from a potentially extensive set of options, along with the determination of the associated power cost.
@@
Game theory has long been an important tool in the design and analysis of communication systems [7], and recent years have seen the use of auctions linked to WPT in conjunction with energy markets [8]. These works, however, apply auctions in the fiscal domain, using money as the payment medium. In contrast, we propose the use of wireless power itself as an RF-native “currency”, thus eliminating the need for complex configuration and integration between the physical wireless system and financial infrastructure.
@@
  • We validate our theoretical findings through extensive simulation experiments, with communication and WPT parameter values taken from real-world experiments reported in the literature. Our results demonstrate a 91% reduction in outage probability with as few as 4 relay candidates compared to the non-cooperative alternative, and as much as 48% source power savings compared to a baseline approach.
@@ II-A Communication System Model
Based on [2, 10], we use a reference path loss channel model with lognormal small scale fading, and we account for large scale fading by considering different sets of channel parameters for LoS and NLoS cases. In the dB scale, this model is parameterized by a path loss intercept, $H_{\text{PL},0,\text{dB}}^{c}$, a path loss exponent, $n_{\text{PL}}^{c}$, and fading power, $(\sigma_{\text{SS, dB}}^{c})^{2}$, where $c\in\{\text{L},\,\text{NL}\}$ indicates the corresponding channel parameters for the LoS and NLoS cases. We assume that $H_{i}$ and $H_{si}$ are LoS channels, as depicted in Fig. 1, as otherwise, the candidate would almost certainly not be useful as a relay. We thus express these channels as
@@
where $d_{s}=||q_{s}||_{2}^{2}$ is the distance between the source and the AP, and $H_{\sigma_{\text{SS,dB}}^{\text{NL}},s}\sim\mathcal{N}(0,\,(\sigma_{\text{SS,dB}}^{\text{NL}})^{2})$. The small scale fading is taken to be independent across all channels [10].
@@ II-B Wireless Power Transfer
In order to achieve the utility transfer necessary to enable candidate participation in the aforementioned scenario, we propose leveraging wireless power transfer from the source to the relay candidates. More specifically, the source uses simultaneous wireless information and power transfer (SWIPT), so that the total source transmit power is given by $P^{\text{tot}}=P_{si}+P_{\text{WPT}}$, where $P_{\text{WPT}}$ is the power allocated for WPT. The candidate then uses a dynamic power-splitting setup [11], with the power splitting ratio given by $P_{\text{WPT}}/P^{\text{tot}}$. The power harvested by a candidate relay $i$ is then given by $P_{i}^{\text{harv}}=\alpha P_{\text{WPT}}A_{r}H_{si}$ [12], with $\alpha\in[0,1]$ denoting the efficiency of the power harvesting circuitry, and $A_{r}$ denoting the effective aperture area of the receiving antenna (assumed to be identical for all candidates).
@@ III-B Formal Problem Statement
The interpretation of the auction result in the context of the underlying communication system requires brief consideration. In particular, when $\rho(v)=0$, the source “wins” the auction (no suitable candidate is found), and if, additionally, $P_{s}\leq P_{\text{max}}$, the source communicates directly with the AP with power $x(v)=P_{s}$. However, if $P_{s}>P_{\text{max}}$, the source will be unable to connect to the AP, and $x(v)=P_{\text{max}}$ is no longer representative of the transmit power (which is 00), as no communication occurs. On the other hand, any mechanism that does not result in communication when feasible can be shown to be sub-optimal within the auction framework [^3] A simple proof proceeds by contradiction but is omitted for brevity.. Further, when communication is infeasible, i.e., if $P_{\text{max}}<\min\bigl{(}\{P_{s}\}\cup\{P_{si}^{\text{min}}\}_{i\in N}\bigr{)}$, all mechanisms that respect individual rationality yield the same result, $x(v)=P_{\text{max}}$, so that the optimization is effectively over only scenarios where connectivity is achievable. Thus, optimality within the auction framework corresponds to a protocol in the communication system which (1) always results in connectivity when feasible and (2) minimizes the expectation of source’s transmit power conditioned on the event that connectivity is possible.
@@ IV-A Buyer-Optimal Reverse Auctions
We note that the reverse auction may be transformed into a forward auction with valuations $\{-v_{i}\}_{i=0}^{n}$ and payments $-x(v)$. The result then follows from [9, Lemma 3]. ∎
@@ IV-B Optimal WPT Auction
Figure 2: Sample simulation results: (a) Consistent reduction in outage probability with a Myerson auction-based relay selection, incentivized by WPT for (left) various values of WPT circuitry efficiency, $\alpha$ with $\Gamma=1$ and (right) several scales of small-sacle fading variance, $\Gamma$ with $\alpha=0.2$. (b) Higher relative source power savings of the Myerson mechanism with increasing number of relay candidates (c) Variation in harvested power with WPT efficiency ($\alpha$) as the number of relay candidates grow, with an overall reduction due to higher source power savings.
@@
$2+\frac{1-\Phi(z)}{\phi(z)}[\sigma_{\text{SS}}^{\text{L}}-z]>2-\frac{1-\Phi(z)}{\phi(z)}z>1,\,\forall z>0.\vspace{-0.1in}$
@@ V Numerical Examples
To verify the validity of our proposed approach, we run a number of simulation experiments with system parameters as reported from experimental studies. In particular, we consider the scenario depicted in Fig. 1, in which a source seeks to connect with an access point with which it has a NLoS channel. We randomly place up to 10 relay candidates with LoS channels to both the AP and the source. We then follow the auction protocol described in Section IV to select a candidate relay (if any) and determine the transmit power of the source. Throughout, we assume all links communicate using 8-QAM modulation, with a target bit error rate (BER) of $1\mathrm{e}{-6}$, so that the target SNR in dB is given by $\gamma_{\text{th,dB}}=33.18$ [16]. Communication noise is modeled as AWGN with power $\sigma^{2}=-75\,$dBm, and the maximum transmit power of the source is 100100100\,100mW. The channel parameters are given as follows: $H_{\text{PL},0,dB}^{\text{L}}=0$, $H_{\text{PL},0,dB}^{\text{NL}}=-25$, $n_{\text{PL}}^{\text{L}}=2.5$, $n_{\text{PL}}^{\text{NL}}=5.76$, $\sigma_{SS,dB}^{\text{L}}=8.66$, and $\sigma_{SS,dB}^{\text{NL}}=9.06$ [2, Table VI][17]. For WPT, we take $\alpha=0.2$ and $A_{r}=1\,\text{cm}^{2}$. For several experiments, we vary both $\sigma_{SS,dB}^{\text{L}}$ and $\sigma_{SS,dB}^{\text{NL}}$ by scaling the given values with a factor, $\Gamma\in\{0.2,0.6,1.0,1.4\}$. Each reported average is taken over a set of 10,000 simulations with randomly sampled channel fading, and, unless stated otherwise, random placement of the candidates.
@@ V-A Outage Probability
We first study the impact of our protocol on the outage probability of the source. In the absence of relay candidates, the source experiences an outage probability of 969696\,96%. Fig. 2 (a) plots the outage probability of the source for different values of $n$, $\alpha$ and fading scale factors, $\Gamma$. We see that the presence of even a single relay candidate significantly reduces the outage probability to as little as 454545\,45%, and this probability drops to near 000\,for higher values of $n$. Furthermore, increasing either power efficiency, $\alpha$, or fading factor, $\Gamma$, reduces outage probability.
@@ V-B Source Transmit Power
We next examine the transmit power used by the source as determined by the auction protocol. To gauge energy efficiency, we compare the transmit power, $P^{tot}$, of the source under the Myerson mechanism to the perfect-information transmit power, $P^{\text{lb}}$ (described in Section III-B), which is a lower bound on $P^{\text{tot}}$. As an upper bound, we also compare against a standard sealed-bid, second-price auction (Vickrey auction), in which the source selects the candidate relay with the lowest valuation but transmits with a power equal to the second-lowest valuation [18]. Fig. 2 (b) shows significant power savings when using the proposed buyer-optimal Myerson auction protocol compared to a standard Vickrey auction, particularly when fewer relay candidates are present. Furthermore, as the number of candidates increase, our proposed Myerson auction solution approaches the perfect-information lower bound. We also note that the performance of the Vickrey and Myerson auctions become comparable when the number of candidates is large, suggesting that in scenarios where candidates and sources have less information about each other’s channels, the Vickrey auction may produce acceptable results given a sufficiently high density of UEs.
@@ V-C Harvested Transmission Power
While energy harvesting is not the chief objective of our protocol, it plays an important role in motivating the candidates to cooperate. Fig. 2 (c) shows the average amount of power harvested by candidates for various efficiency factors, $\alpha$. As expected, more energy is harvested as the WPT efficiency is increased, and we see that the candidates receive small but significant boosts on the order of tens of $\mu$W. However, as the number of candidates grow, competition allows the source to reduce the transmit power, $P^{\text{tot}}$, which in turn reduces the power available for harvesting.

### v2

@@ Abstract
This paper addresses the challenge of incentivizing energy-constrained, non-cooperative user equipment (UE) to serve as cooperative relays. We consider a source UE with a non-line-of-sight channel to an access point (AP), where direct communication may be infeasible or may necessitate a substantial transmit power. Other UEs in the vicinity are viewed as relay candidates, and our aim is to enable energy-efficient connectivity for the source, while accounting for the self-interested behavior and private channel state information of these candidates, by allowing the source to “pay” the candidates via wireless power transfer (WPT). We propose a cooperation-inducing protocol, inspired by Myerson auction theory, which ensures that candidates truthfully report power requirements while minimizing the expected power used by the source. Through rigorous analysis, we establish the regularity of valuations for lognormal fading channels, which allows for the efficient determination of the optimal source transmit power. Extensive simulation experiments, employing real-world communication and WPT parameters, validate our theoretical framework. Our results demonstrate over 71% reduction in outage probability with as few as 4 relay candidates, compared to the non-cooperative scenario, and as much as 70% source power savings compared to a baseline approach, highlighting the efficacy of our proposed methodology.
@@ I Introduction
The forthcoming 6G wireless standard promises to revolutionize wireless communication by delivering significant improvements in reliability, data rates, and energy efficiency [1]. Achieving the desired technical performance depends on several emerging technologies and paradigms, which in turn bring their own challenges. For instance, high frequency communication (e.g., mmWave, THz) enables vastly increased capacity, but suffers from acute shadowing and penetration losses [2]. At the same time, given the ever-increasing number of connected devices, there is a growing need for scalable system design, with reduced reliance on centralized coordination mechanisms, to produce the desired performance.
This paper proposes the use of wireless power transfer (WPT) [3] to induce cooperative behavior among non-cooperative communication devices, with the end goal of enabling connectivity and/or reducing energy consumption. More specifically, we consider a setting in which a large-scale blockage (e.g., a wall) obstructs the line-of-sight (LoS) link between a user equipment (UE) and an access point (AP), so that direct communication is challenging. To reduce the power required to connect to the AP, the blocked user may offer to send a payment of power (via WPT) to a more advantageously placed UE, in exchange for relay services. The core problem we address in this paper is the identification of the optimal relay candidate from a potentially extensive set of options, along with the determination of the associated power cost.
@@
Game theory has long been an important tool in the design and analysis of communication systems [7, 8], and recent years have seen the use of auctions linked to WPT in conjunction with energy markets [9]. These works, however, apply auctions in the fiscal domain, using money as the payment medium. In contrast, we propose the use of wireless power itself as an RF-native “currency”, thus eliminating the need for complex configuration and integration between the physical wireless system and financial infrastructure.
@@
  • We validate our theoretical findings through extensive simulation experiments, with communication and WPT parameter values taken from real-world experiments reported in the literature. Our results demonstrate a 71% reduction in outage probability with as few as 4 relay candidates compared to the non-cooperative alternative, and as much as 48% source power savings compared to a baseline approach.
@@ II-A Communication System Model
Based on [2, 11], we use a reference path loss channel model with lognormal small-scale fading, and we account for large scale fading by considering different sets of channel parameters for LoS and NLoS cases. In the dB scale, this model is parameterized by a path loss intercept, $H_{\text{PL},0,\text{dB}}^{c}$, a path loss exponent, $n_{\text{PL}}^{c}$, and fading power, $(\sigma_{\text{SS, dB}}^{c})^{2}$, where $c\in\{\text{L},\,\text{NL}\}$ indicates the corresponding channel parameters for the LoS and NLoS cases. We assume that $H_{i}$ and $H_{si}$ are LoS channels, as depicted in Fig. 1, as otherwise, the candidate would almost certainly not be useful as a relay. We thus express these channels as
@@
where $d_{s}=||q_{s}||_{2}^{2}$ is the distance between the source and the AP, and $H_{\sigma_{\text{SS,dB}}^{\text{NL}},s}\sim\mathcal{N}(0,\,(\sigma_{\text{SS,dB}}^{\text{NL}})^{2})$. The small-scale fading is taken to be independent across all channels [11].
@@ II-B Wireless Power Transfer
To achieve the utility transfer necessary to enable candidate participation in the aforementioned scenario, we propose leveraging wireless power transfer from the source to the relay candidates. More specifically, the source uses simultaneous wireless information and power transfer (SWIPT), so that the total source transmit power is given by $P^{\text{tot}}=P_{si}+P_{\text{WPT}}$, where $P_{\text{WPT}}$ is the power allocated for WPT. The candidate then uses a dynamic power-splitting setup [12], with the power splitting ratio given by $P_{\text{WPT}}/P^{\text{tot}}$. The power harvested by a candidate relay $i$ is then given by $P_{i}^{\text{harv}}=\alpha P_{\text{WPT}}A_{r}H_{si}$ [13], with $\alpha\in[0,1]$ denoting the efficiency of the power harvesting circuitry, and $A_{r}$ denoting the effective aperture area of the receiving antenna (assumed to be identical for all candidates).
@@ IV-A Buyer-Optimal Reverse Auctions
We note that the reverse auction may be transformed into a forward auction with valuations $\{-v_{i}\}_{i=0}^{n}$ and payments $-x(v)$. The result then follows from [10, Lemma 3]. ∎
@@ IV-B Optimal WPT Auction
Figure 2: Sample simulation results: (a) Consistent reduction in outage probability with a Myerson auction-based relay selection, incentivized by WPT for (left) various values of WPT circuitry efficiency, $\alpha$ with $\Gamma=1$ and (right) several scales of small-scale fading variance, $\Gamma$ with $\alpha=0.3$. (b) Higher relative source power savings of the Myerson mechanism with increasing number of relay candidates (c) Variation in harvested power with WPT efficiency ($\alpha$) as the number of relay candidates grows, with an overall reduction due to higher source power savings. See color PDF for optimal viewing.
@@
$2+\frac{1-\Phi(z)}{\phi(z)}[\sigma_{\text{SS}}^{\text{L}}-z]>2-\frac{1-\Phi(z)}{\phi(z)}z>1,\,\forall z>0.\qed\vspace{-0.1in}$
@@
The interpretation of the auction result in the context of the underlying communication system requires brief consideration. When $\rho(v)=0$, the source “wins” the auction (no suitable candidate is found), and if, additionally, $P_{s}\leq P_{\text{max}}$, the source communicates directly with the AP with power $x(v)=P_{s}$. However, if $P_{s}>P_{\text{max}}$, the source will be unable to connect to the AP, and $x(v)=P_{\text{max}}$ is no longer representative of the transmit power (which is 00), as no communication occurs. In particular, communication will occur under the perfect info baseline but will not occur under the Myerson auction protocol if the following three conditions are met: the source cannot communicate directly ($P_{s}>P_{\text{max}}$), the source wins the auction ($c_{i}(v_{i})>P_{\text{max}}\forall i\in N$), and there exists at least one candidate, $i$, such that $P_{si}^{\text{min}}\leq P_{\text{max}}$. However, as we will see the experimental results of the Section V, the Myerson auction still greatly increases the likelihood of successful communication.
@@ V Numerical Examples
To verify the validity of our proposed approach, we run several simulation experiments with system parameters as reported from experimental studies. We consider the scenario depicted in Fig. 1, in which a source seeks to connect with an access point with which it has a NLoS channel. We randomly place up to 10 relay candidates with LoS channels to both the AP and the source. We then follow the auction protocol described in Section IV to select a candidate relay (if any) and determine the transmit power of the source. Throughout, we assume all links communicate using 8-QAM modulation, with a target bit error rate (BER) of $1\mathrm{e}{-6}$, so that the target SNR in dB is given by $\gamma_{\text{th,dB}}=33.18$ [17]. Communication noise is modeled as AWGN with power $\sigma^{2}=-75\,$dBm, and the maximum transmit power of the source is 100100100\,100mW. The channel parameters are given as follows: $H_{\text{PL},0,dB}^{\text{L}}=0$, $H_{\text{PL},0,dB}^{\text{NL}}=-25$, $n_{\text{PL}}^{\text{L}}=2.5$, $n_{\text{PL}}^{\text{NL}}=5.76$, $\sigma_{SS,dB}^{\text{L}}=8.66$, and $\sigma_{SS,dB}^{\text{NL}}=9.06$ [2, Table VI][18]. For WPT, we take $A_{r}=1\,\text{cm}^{2}$ and consider $\alpha\in\{0.1,0.2,0.3,0.4\}$. For several experiments, we vary both $\sigma_{SS,dB}^{\text{L}}$ and $\sigma_{SS,dB}^{\text{NL}}$ by scaling the given values with a factor, $\Gamma\in\{0.2,0.6,1.0,1.4\}$. Each reported average is taken over a set of 10,000 simulations with randomly sampled channel fading, and, unless stated otherwise, random placement of the candidates.
@@ V-A Outage Probability
We first study the impact of our protocol on the outage probability of the source. Fig. 2 (a) plots the outage probability of the source for different values of $n$, $\alpha$ and fading scale factors, $\Gamma$. In the absence of relay candidates, the source experiences an outage probability of between $1$ and $0.89$, depending on the value of $\Gamma$, but we see that the presence of as few as four relay candidates, the outage probability drops below $0.2$. This probability drops to near 00 for higher values of $n$. Furthermore, increasing either power efficiency, $\alpha$, or fading factor, $\Gamma$, reduces outage probability.
@@ V-B Source Transmit Power
We next examine the transmit power used by the source as determined by the auction protocol. To gauge energy efficiency, we compare the transmit power, $P^{\text{tot}}$, of the source under the Myerson mechanism to the perfect-information transmit power, $P^{\text{lb}}$ (described in Section III-B), which is a lower bound on $P^{\text{tot}}$. As an upper bound, we also compare against a standard sealed-bid, second-price auction (Vickrey auction), in which the source selects the candidate relay with the lowest valuation but transmits with a power equal to the second-lowest valuation [19]. Fig. 2 (b) shows significant power savings when using the proposed buyer-optimal Myerson auction protocol compared to a standard Vickrey auction, particularly when fewer relay candidates are present. Furthermore, as the number of candidates increases, our proposed Myerson auction solution approaches the perfect-information lower bound. We also note that the performance of the Vickrey and Myerson auctions become comparable when the number of candidates is large, suggesting that in scenarios where candidates and sources have less information about each other’s channels, the Vickrey auction may produce acceptable results given a sufficiently high density of UEs.
@@ V-C Harvested Transmission Power
While energy harvesting is not the chief objective of our protocol, it plays an important role in motivating the candidates to cooperate. Fig. 2 (c) shows the average amount of power harvested by candidates for various efficiency factors, $\alpha$. As expected, more energy is harvested as the WPT efficiency is increased, and we see that the candidates receive small but significant boosts, on the order of tens of $\mu$W. However, as the number of candidates grows, competition allows the source to reduce the transmit power, $P^{\text{tot}}$, which in turn reduces the power available for harvesting.